### v2

# Investigating The Impact of Quantization on Factual Knowledge Recall

Abstract
Quantization methods are widely used to accelerate inference and streamline the deployment of large language models (LLMs). Although quantization’s effects on various LLM capabilities have been extensively studied, one critical area remains underexplored: factual knowledge recall (FKR), the process by which LLMs access stored knowledge. To this end, we conduct comprehensive experiments using three common quantization techniques at distinct bit widths, in conjunction with interpretability-driven analyses on two tasks, knowledge memorization and latent multi-hop reasoning.We show that quantization typically results in information loss within LLMs, consequently diminishing their capacity for FKR. This effect is particularly amplified in smaller models within the same architectural families. However, models quantized at reduced bit precision do not consistently exhibit inferior performance and occasionally quantization may even enhance model FKR. We find that BitSandBytes demonstrates highest preservation of the original full-precision model’s FKR. Despite variability across models and methods, quantization causes modest performance degradation and remains an effective compression strategy.
[Uncaptioned image] Through a Compressed Lens:
Investigating The Impact of Quantization on Factual Knowledge Recall
Qianli Wang1,2    Mingyang Wang4,5,6    Nils Feldhus1,2,8    Simon Ostermann2,3,7 Yuan Cao9   Hinrich Schütze4,6   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin  2German Research Center for Artificial Intelligence (DFKI) 3Saarland Informatics Campus  4LMU Munich  5Bosch Center for Artificial Intelligence (BCAI) 6Munich Center for Machine Learning (MCML)  7Centre for European Research in Trusted AI (CERTAIN) 8BIFOLD – Berlin Institute for the Foundations of Learning and Data  9Technical University of Munich Correspondence: qianli.wang@tu-berlin.de
1 Introduction
The shift towards LLMs has created strong demand for efficient inference and accessible deployment. In response, numerous quantization techniques have been developed dettmers-etal-2022-8bits; frantar-etal-2023-optq; xiao-etal-2023-smooth; lin-etal-2024-awq. By reducing the precision of a model’s parameters, quantization allows us to decrease model size while mostly preserving its performance gray-etal-1998-quantization. Although the effects of quantization in LLMs have been evaluated across various aspects, e.g., multilinguality, bias, fairness, and trustworthiness marchisio-etal-2024-quantization; goncalves-strubell-2023-understanding; ramesh-etal-2023-comparative; hong-etal-2024-trust; liu-etal-2024-emergent, its impact on LLMs’ ability to recall factual knowledge from pretrained memory remains underexplored and has yet to be comprehensively characterized.
Figure 1: The effect of quantization on factual knowledge recall through knowledge memorization analysis and latent multi-hop reasoning analysis.
To this end, we present a comprehensive study of quantization’s impact on factual knowledge recall (FKR) in LLMs. Instead of solely observing the performance degradation, we perform interpretability-driven analyses on knowledge memorization and latent multi-hop reasoning3.1) tasks, to scrutinize behavior across neuron, layer, and model levels and evaluate three LLMs across three widely used model quantization techniques at different bit widths (Figure 1). We reveal that quantization typically impairs FKR, which is particularly acute for smaller models within a given model family. LLMs quantized to a lower bit precision do not consistently underperform those with higher precision. In some cases, quantization can even paradoxically enhance factual recall. Our findings indicate that BitSandBytes is most effective at preserving the model’s FKR. We observed no significant FKR degradation from quantization that would compromise model compression effectiveness, though effects vary by model and technique.
2 Background and Related Work
Quantization.
Quantization techniques compress LLMs by converting model weights, activations, or the KV cache into lower-precision data types zhu-2024-surveymodelcompressionlarge. These techniques can be broadly categorized into two types: quantization-aware training (QAT) and post-training quantization (PTQ). QAT requires retraining to mitigate errors introduced by quantization, whereas PTQ facilitates the direct application of a quantized model during inference. In this paper, we primarily evaluate the impact of weight-only quantization (§3.3) on FKR (§4), which eliminates the need for retraining to address errors resulting from quantization.
Impact of Quantization. Recent work has extensively examined the impact of quantization on various capabilities of LLMs. marchisio-etal-2024-quantization conduct a thorough analysis of quantized multilingual LLMs, focusing on performance degradation across languages. goncalves-strubell-2023-understanding; kirsten-etal-2024-bias explore the emergence of bias in the outputs generated by quantized models. liu-etal-2024-emergent find that in-context learning ability gradually declines in heavily quantized LLMs. jin-etal-2024-comprehensive observe that models with 4-bit quantization can still retain the alignment ability. singh2025interpretingeffectsquantizationllms investigate the impact of quantization on model calibration. In our work, we explicitly explore how quantization affects FKR.
3 Experimental Setup
We examine two representative methods (§3.1) to evaluate the impact of quantization on FKR. Specifically, we compare full-precision LLMs (§3.4) with LLMs quantized using different techniques and bit configurations (§3.3) across two datasets (§3.2).
3.1 Methods
Knowledge Memorization Analysis.
We investigate the memorization of a model’s factual knowledge, identifying the reasons behind potential factual forgetting namburi-etal-2023-cost. By leveraging the theory of knowledge neurons, which suggests that specific neurons in LLMs are responsible for storing specific pieces of knowledge, we explore how quantization alters the storage and retrieval processes within these neurons.
Latent Multi-hop Reasoning Analysis.
We adapt the methodology from yang-2024-latent-multi-hop-reasoning for inspecting latent multi-hop reasoning errors. Specifically, the analysis tests whether full-precision and quantized LLMs employ similar latent reasoning pathways and internal knowledge recall mechanisms to answer complex factual queries.
3.2 Datasets
Our study employs two widely recognized datasets [^1] Dataset examples are detailed in Appendix A. for evaluating FKR with selected interpretability methods (§3.1).
LRE hernandez-2024-linearity is a knowledge probing dataset consisting of knowledge triplets $(s,r,o)$, structured in a one-hop setting, where $s$, $r$, and $o$ represents the subject, relation, and object, respectively.
TwoHop-Fact yang-2024-latent-multi-hop-reasoning is a dataset consisting of pairs of prompts: two-hop prompts ($\tau_{2H}$) for compositional queries, representing fact composition queries in the form $((e_{1},r_{1},e_{2}),(e_{2},r_{2},e_{3}))$, where $r_{1}$ and $r_{2}$ are relations and $e_{i}$ denotes an entity; and one-hop prompts ($\tau_{1H}$) for subqueries (Figure 4). $e_{1}$ and $e_{2}$ (in the second triplet) are subjects, while $e_{2}$ (in the first triplet) and $e_{3}$ objects. $e_{2}$ serves as a bridge entity, linking the two triplet to form a coherent two-hop reasoning chain.
3.3 Quantization Techniques
Building on the prior discussion (§2), we identify three commonly used PTQ techniques applied to the selected LLMs (§3.4) in our experiments:
  • GPTQ frantar-etal-2023-optq uses a second-order, Hessian-based optimization to quantize weights post-training with minimal accuracy loss;
  • AWQ lin-etal-2024-awq enhances weight quantization by handling activation outliers to preserve model accuracy at low bit-widths;
  • Integer quantization dettmers-etal-2022-8bits implemented by BitsAndBytes (bib4 and bib8) enables fast and memory-efficient inference by using optimized low-bit kernels.
3.4 Models
We evaluate Llama3-8B llama3modelcard, Qwen2.5-7B and Qwen2.5-14B qwen2024qwen25technicalreport using two interpretability methods (§3.1). These models are selected due to the availability of quantized versions for each [^2] Detailed information for each model, including links to their respective quantized versions, is provided in Table 1., ensuring the reproducibility of our results.
(a) Top neuron distribution (Qwen2.5-7B)
(b) Top neuron distribution (Llama3-8B)
(c) Attention (Qwen2.5-7B): Landmark on continent
(d) FFN (Qwen2.5-7B): Landmark on continent
Figure 2: Top: Distribution of high-contributing neurons across layers, showing the average number of top-300 neurons per layer for Qwen2.5-7B (left) and Llama3-8B (right). Bottom: Layer-wise drop in neuron contribution scores across quantization methods for the landmark on continent relation in Qwen2.5-7B, comparing attention sublayers (left) and feed-forward sublayers (right).
4 Evaluation Setup
Knowledge Memorization Analysis.
To assess the effects of quantization on FKR, we first evaluate factual recall accuracy using LRE. A comparison of results before and after quantization reveals the degree of knowledge forgetting. We then employ a knowledge attribution method (yu-ananiadou-2024-neuron) to trace the information loss back to specific layers and neurons, following the knowledge neuron theory. This combined analysis uncovers how quantization impacts the internal mechanisms responsible for storing and retrieving information.
Latent Multi-hop Reasoning Analysis (LMHR).
[…]
5 Results
5.1 Knowledge Memorization Analysis
Table 2 shows that quantization introduces varying degrees of accuracy drop depending on both the model size and quantization method. Notably, the accuracy drop is more obvious in smaller models within the same model family. Additionally, the accuracy of AWQ, GPTQ4, and bib4 quantized models consistently drop across all model sizes, whereas GPTQ8 and bib8 models retain performance comparable to the full-precision models. Furthermore, relations where full-precision performance has not saturated tend to exhibit more severe factual knowledge recall degradation (Table 4), indicating that such knowledge is more fragile under quantization.
Neuron-level Trends. Following the neuron-level method of yu-ananiadou-2024-neuron, we assign each neuron in the model a contribution score equal to the increase in log-probability of the correct answer token that the neuron induces. Comparing these scores before and after quantization allows us to quantify how much individual neurons’ influence changes. We analyze four relations that suffer a notable accuracy drop: landmark on continent, person father, person mother, and person sport position.
We track the top-300 feed-forward neurons in full-precision model, set $\tau$ as their minimum contribution score, and count how many neurons in quantized models exceed $\tau$. As shown in Figure 2 and 2, the per-layer counts decrease after quantization for both models, with the most obvious reductions in the last layers. Complete results across relations are provided in Figures 7 and 8 in App. C.1.
Layer-wise Trends. Figure 2 depicts the drop in aggregate contribution scores for attention (left) and feed-forward (right) sub-layers. All quantization methods exhibit a pronounced decline in the final two layers on Qwen2.5-7B, this trend is consistent on all relations we investigate, as shown in Figure 5 in App C.1. While on Llama3-8B (see Figure 6), the decline occurs in middle-to-late layers, with an increase in the final layers, indicating that information loss patterns vary across model architectures. This divergence may be ascribable to the different ways in which factual knowledge is stored across model families choe2025autoregressivetransformersrememberfacts.
Collectively, both layer-level and neuron-level analyses reveal that quantization primarily affects the network’s last layers. These findings confirm that quantization degrades the decisive information stored in these late layers, accounting for the factual knowledge recall degradation reported in Table 2.333The complete analysis of both models across different relations is provided in Appendix C.1.
(a) Qwen2.5-7B
(b) Qwen2.5-14B
(c) Llama3-8B
Figure 3: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the GPTQ8 quantized and full-precision models, evaluated across all layers.
5.2 Latent Multi-hop Reasoning Analysis
Quantization affects the first-hop Reasoning the most. Table 3 reveals that quantization substantially affects the first hop $r_{1}(e_{1})$, by as much as $30.08\%$, while its impact on the second hop is minimal, with an average of degradation of only $4.25\%$. The $r_{2}(r_{1}(e_{1}))$ deterioration due to quantization is strongly correlated with the ability to correctly predict the bridge entities $r_{1}(e_{1})$, as indicated by a Spearman’s correlation of $0.93$. Nevertheless, the FKR deterioration is not dramatic and considerably acceptable (e.g., Qwen2.5-7B shows a minor average deterioration of 0.77%).
Quantization Effects are not consistent. Figure 3 illustrates that quantization effects on FKR are largely unpredictable. The quantization effect is heterogeneous across layers and variable across models, particularly across architectures, given that facts are stored in various ways across Llama3 and Qwen2.5 models choe2025autoregressivetransformersrememberfacts. Besides, different quantization approaches often affect the model FKR in different manners (Figure 12). Nevertheless, for a given quantization method with different bit width, the layer-wise impact remains broadly similar. Surprisingly, quantization can occasionally even improve FKR (Figure 3). This effect may be attributed to the regularization effect park-etal-2022-robust or quantization-induced noise li2024investigating, which can inadvertently enhance the model’s capability to recall factual knowledge.
bib largely preserve the factual knowledge recall capability. For Qwen2.5-7B, quantization consistently reduce knowledge recall, with the largest degradations under GPTQ and AWQ (Figure 12). By contrast, bib4/8 degrades one-hop reasoning $r_{1}(e_{1})$ and $r_{2}(e_{2})$, as reflected by modest shifts in EntRec and CnstScore relative to the full-precision model, but leaves two-hop reasoning largely unaffected. For Qwen2.5-14B, bib4/8 likewise outperforms other quantization methods, best preserving and occasionally improving FKR (Figure 14). Moreover, despite its greater layer depth than Qwen2.5-7B, Qwen2.5-14B exhibits a similar layer-wise pattern of bib4/8 quantization effects (Figure 12(e), 14(e)), whereas other quantization methods do not. For Llama3-8B, the effect of different quantization methods are alike, i.e., EntRec and CnstScore of quantized model are lower than full-precision model in shallow layers, but becomes much higher in deeper layers (Figure 15). Among the three LLMs evaluated, Llama3-8B is least affected by quantization in terms of FKR.
6 Conclusion
In this work, we examined the impact of quantization on factual knowledge recall. In general, quantization leads to information loss within the model, which in turn degrades factual knowledge recall; this trend becomes more pronounced for smaller models within each model family. Moreover, LLMs quantized with lower bit precision do not invariably perform worse than those with higher bit precision. Quantization can occasionally even improve factual knowledge recall. While quantization effects vary by model and technique, we observed no performance degradation severe enough to compromise its viability as a compression strategy.
Limitations
Our experimental work is confined to English-language datasets. Consequently, the effectiveness of our experiments in other languages may not be comparable and multilingual factual knowledge recall may be simultaneously affected by the degradation of multilingual capabilities due to quantization marchisio-etal-2024-quantization. Extending experiments to the multilingual settings is considered for future work.
We restrict our experiments to Llama3-8B, Qwen2.5-7B and Qwen2.5-14B for computational feasibility: our gradient-based analyses exceed the available GPU memory for larger models. Between the model families (Qwen, Llama, Mistral, DeepSeek), there are lots of architectural equivalences and similarities, e.g., the same attention (grouped-query attention), position embeddings (RoPE), normalization (RMSNorm) or FFN activation (SwiGLU). We argue that, based on our comprehensive experiments, our results are considerably robust and generalizable given the similar architectures compared to other model families.
In our experiments, we extensively compare full-precision models with different quantized versions in 4-bit and 8-bit formats. Lower-bit quantization, such as 1-bit or 2-bit, is not included in our study.
Moreover, the scope of our experiments is limited to post-training quantization techniques. Investigating the impact of weight-activation quantization, KV cache compression, or quantization-aware training techniques on factual knowledge recall is counted as future work.
Appendix A Dataset Information
LRE
An exemplary query consists of $s\leftarrow\text{"{the company that created Visual Basic}"}$ and $o\leftarrow\text{"{The current CEO of}"}$. A correct answer (in this example: Satya Nadella) by the explained model is the criterion by which the LRE data is filtered.
TwoHop-Fact
Figure 4 illustrates an example from the TwoHop-Fact dataset. Based on the input text, $(e_{1},r_{1},e_{2})$ corresponds (Superstition, singer, Stevie Wonder), while $(e_{2},r_{2},e_{3})$ represents (Stevie Wonder, mother, Lula).
TwoHop-Fact (Multi-hop Reasoning) One-hop prompts: ($p_{1}$): The mother of Stevie Wonder is Lula. ($p_{2}$): The singer of ‘Superstition’ is Stevie Wonder. Two-hop prompt: ($p$): The mother of the singer of ‘Superstition’ is Answer: Lura
Figure 4: An example from the TwoHop-Fact dataset.
Appendix B Models & Inference Time
Name Citation Size Precision Link
Llama3 llama3modelcard 8B full https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct
Llama3 llama3modelcard 8B GPTQ4 https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-GPTQ
Llama3 llama3modelcard 8B AWQ https://huggingface.co/TechxGenus/Meta-Llama-3-8B-Instruct-AWQ
Qwen2.5 qwen2024qwen25technicalreport 7B Full https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Qwen2.5 qwen2024qwen25technicalreport 7B AWQ https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-AWQ
Qwen2.5 qwen2024qwen25technicalreport 7B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int4
Qwen2.5 qwen2024qwen25technicalreport 7B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-7B-Instruct-GPTQ-Int8
Qwen2.5 qwen2024qwen25technicalreport 14B Full https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5 qwen2024qwen25technicalreport 14B AWQ https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-AWQ
Qwen2.5 qwen2024qwen25technicalreport 14B GPTQ4 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int4
Qwen2.5 qwen2024qwen25technicalreport 14B GPTQ8 https://huggingface.co/Qwen/Qwen2.5-14B-Instruct-GPTQ-Int8
Table 1: Detailed information about used LLMs in our experiments.
Table 1 presents details of the all LLMs used in our experiments (§3.4), including model sizes, quantization approaches and corresponding URLs from the Hugging Face Hub. All models were directly obtained from the Hugging Face repository. All experiments were conducted using A100 or H100 GPUs. Neuron-level and layer-level attribution can be completed within 10 hours, while LMHR experiments take 30 hours averagely.
Appendix C Additional Experiments
C.1 Knowledge Memorization Analysis
Model Method Accuracy % $\uparrow$ # Correct
Qwen2.5-7B full 63.25 6133
bib4 60.72 5887
bib8 63.01 6109
gptq4 60.10 5827
gptq8 63.22 6130
awq 60.60 5876
Qwen2.5-14B full 73.08 7086
bib4 70.33 6819
bib8 73.06 7084
gptq4 25.20 2443
gptq8 73.03 7081
awq 70.61 6846
Llama3-8B full 77.62 7526
bib4 72.19 7000
bib8 76.95 7461
gptq8 71.39 6922
awq 71.83 6965
Table 2: Knowledge recall accuracy results (%) and number of correct predictions (out of 9696 queries) on the LRE dataset for Qwen2.5-{7B,14B} and Llama3-8B models across different quantization methods.
Table 2 illustrates knowledge recall accuracy results on the LRE dataset for Qwen2.5-{7B,14B} and Llama3-8B models across different quantization methods.
Knowledge Recall Accuracy.
In Table 4, and 5, we provide the per-relation factual recall accuracy on Qwen2.5-7B, Qwen2.5-14B, and Llama3-8B models, respectively. As discussion in Section 5.1, we observe that the accuracy drop is more severe in relations where the original performance has not yet saturated.
Neuron Attribution Analysis.
Here we present the neuron-level knowledge attribution results on relations person father, person mother, person sport position, complementing our results in Section 5.1. Analysis on layer-wise neuron contribution scores are shown in Figure 5 and Figure 6, the analysis on top-300 neurons are given in Figure 9 and Figure 10.
(a) Attention: Person father
(b) FFN: Person father
(c) Attention: Person mother
(d) FFN: Person mother
(e) Attention: Person sport position
(f) FFN: Person sport position
Figure 5: Contribution score drop across quantization methods on Qwen2.5-7B for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right). Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position. The plots reveal how different quantization methods affect knowledge representation across model layers.
(a) Attention: Person father
(b) FFN: Person father
(c) Attention: Person mother
(d) FFN: Person mother
(e) Attention: Person sport position
(f) FFN: Person sport position
Figure 6: Contribution score drop across quantization methods on Llama3-8B for other relationship types, comparing attention sublayers (left) and feed-forward sublayers (right). Rows show different relationships: (a-b) person father, (c-d) person mother, and (e-f) person sport position. The plots reveal how different quantization methods affect knowledge representation across model layers.
(a) Contribution score comparison
(b) Layer distribution
Figure 7: Analysis of the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods applied to Qwen2.5-7B. (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns. (b) Distribution of high-scoring neurons across layers, showing the number of neurons exceeding the full-precision model’s 300th neuron score threshold in each layer.
(a) Contribution score comparison
(b) Layer distribution
Figure 8: Analysis of the top-300 neurons with highest contribution scores for the landmark on continent relation under different quantization methods applied to Llama3-8B. (a) Average contribution scores of top 300 feed-forward neurons across different quantization methods, showing how each method affects neuron activation patterns. (b) Distribution of high-scoring neurons across layers, showing the number of neurons exceeding the full-precision model’s 300th neuron score threshold in each layer.
Refer to caption Refer to caption
(a) Person father: Contribution scores (b) Person father: Layer distribution
Refer to caption Refer to caption
(c) Person mother: Contribution scores (d) Person mother: Layer distribution
Refer to caption Refer to caption
(e) Person sport position: Contribution scores (f) Person sport position: Layer distribution
Figure 9: Analysis of other relationship types under different quantization methods applied to Qwen2.5-7B. Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods. Right column: Distribution of high-scoring neurons across model layers. Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position. These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
Refer to caption Refer to caption
(a) Person father: Contribution scores (b) Person father: Layer distribution
Refer to caption Refer to caption
(c) Person mother: Contribution scores (d) Person mother: Layer distribution
Refer to caption Refer to caption
(e) Person sport position: Contribution scores (f) Person sport position: Layer distribution
Figure 10: Analysis of other relationship types under different quantization methods applied to Llama3-8B. Left column: Average contribution scores of top 300 feed-forward neurons across quantization methods. Right column: Distribution of high-scoring neurons across model layers. Each row represents a different relationship: (a-b) person father, (c-d) person mother, and (e-f) person sport position. These visualizations reveal both the magnitude of contribution score changes and their distribution across the model architecture when applying different quantization techniques to various types of factual knowledge.
C.2 Latent Multi-hop Reasoning
Model Method $r_{1}(e_{1})$ $\uparrow$ $r_{2}(e_{2})$ $\uparrow$ $r_{2}(r_{1}(e_{1}))$ $\uparrow$
Qwen2.5-7B full 25.03 39.07 20.61
bib4 25.01 39.02 20.61
bib8 25.01 39.02 20.61
gptq4 22.29 37.44 19.59
gptq8 24.76 39.15 20.55
awq 22.07 38.89 18.05
Qwen2.5-14B full 35.23 40.45 24.72
bib4 35.16 40.56 24.76
bib8 35.16 40.56 24.76
gptq4 32.22 45.19 23.29
gptq8 35.16 39.55 24.61
awq 24.69 35.61 21.80
Llama3-8B full 7.79 21.39 4.45
bib4 4.94 27.44 3.14
bib8 5.79 18.27 2.90
gptq8 23.62 40.73 20.94
awq 22.35 37.79 19.56
Table 3: Accuracy (in %) of different models in connecting and traversing implicit knowledge to successfully answer latent multi-hop queries on TwoHop-Fact.
Table 3 shows the latent multi-hop reasoning accuracy comparison between full-precision models and quantized models. Additionally, Figure 12, Figure 14, and Figure 15 display the differences in the entity recall score, consistency score, and accuracy between the AWQ, GPTQ8, GPTQ4, bib8, bib4 quantized and full-precision models of Qwen2.5-7B, Qwen2.5-14B, and Llama3-8B, evaluated across all layers.
(a) AWQ
(b) GPTQ8
(c) GPTQ4
(d) bib8
(e) bib4
Figure 12: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the AWQ, GPTQ8, GPTQ4, bib8, bib4 quantized and full-precision models of Qwen2.5-7B, evaluated across all layers.
(a) AWQ
(b) GPTQ8
(c) GPTQ4
(d) bib8
(e) bib4
Figure 14: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the AWQ, GPTQ8, GPTQ4, bib8, bib4 quantized and full-precision models of Qwen2.5-14B, evaluated across all layers.
(a) AWQ
(b) GPTQ
(c) bib8
(d) bib4
Figure 15: Difference in the entity recall score (EntRec), consistency score (CnstScore), and accuracy between the AWQ, GPTQ, bib8, bib4 quantized and full-precision models of Llama3-8B, evaluated across all layers.
Qwen2.5-7B full bib4 bib8 gptq4 gptq8 awq
city in country 93.00 85.00 93.00 93.00 93.00 93.00
company CEO 49.00 39.00 50.00 45.00 50.00 46.00
company hq 59.00 54.00 59.00 56.00 58.00 59.00
country capital city 100.00 100.00 100.00 100.00 100.00 100.00
country currency 93.00 90.00 93.00 87.00 93.00 97.00
country language 92.00 88.00 92.00 88.00 92.00 92.00
country largest city 100.00 100.00 100.00 100.00 100.00 100.00
food from country 83.00 83.00 83.00 80.00 83.00 80.00
landmark in country 81.00 80.00 81.00 79.00 81.00 80.00
landmark on continent 87.00 84.00 85.00 83.00 87.00 85.00
person father 45.00 39.00 45.00 40.00 45.00 41.00
person lead singer of band 100.00 100.00 100.00 100.00 100.00 100.00
person mother 39.00 32.00 39.00 34.00 38.00 33.00
person native language 90.00 91.00 90.00 88.00 90.00 86.00
person occupation 33.00 34.00 33.00 31.00 33.00 30.00
person plays instrument 44.00 46.00 43.00 42.00 44.00 43.00
person sport position 66.00 63.00 66.00 59.00 66.00 62.00
person university 47.00 46.00 46.00 48.00 46.00 47.00
plays pro sport 79.00 77.00 81.00 75.00 79.00 76.00
pokemon evolution 100.00 89.00 100.00 91.00 100.00 95.00
president birth year 0.00 0.00 0.00 0.00 0.00 0.00
president election year 0.00 0.00 0.00 0.00 0.00 0.00
product by company 80.00 77.00 79.00 77.00 80.00 79.00
star constellation name 85.00 85.00 85.00 85.00 85.00 84.00
superhero archnemesis 34.00 33.00 33.00 33.00 34.00 31.00
superhero person 49.00 50.00 50.00 49.00 50.00 48.00
AVG 63.25 60.72 63.01 60.10 63.22 60.00
Table 4: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-7B across different quantization methods.
Qwen2.5-14B full bib4 bib8 gptq4 gptq8 awq
city in country 96.00 96.00 96.00 81.00 96.00 96.00
company CEO 58.00 54.00 59.00 2.00 58.00 57.00
company hq 68.00 65.00 66.00 10.00 69.00 65.00
country capital city 100.00 100.00 100.00 79.00 100.00 100.00
country currency 90.00 83.00 87.00 77.00 87.00 90.00
country language 96.00 96.00 96.00 92.00 96.00 96.00
country largest city 96.00 100.00 96.00 83.00 96.00 96.00
food from country 80.00 77.00 80.00 50.00 80.00 80.00
landmark in country 87.00 86.00 87.00 31.00 87.00 87.00
landmark on continent 85.00 88.00 84.00 48.00 85.00 84.00
person father 63.00 56.00 62.00 7.00 63.00 57.00
person lead singer of band 100.00 100.00 100.00 29.00 100.00 100.00
person mother 59.00 49.00 59.00 1.00 59.00 53.00
person native language 94.00 94.00 94.00 39.00 93.00 93.00
person occupation 47.00 45.00 49.00 12.00 49.00 45.00
person plays instrument 67.00 61.00 68.00 52.00 66.00 63.00
person sport position 76.00 74.00 76.00 5.00 74.00 71.00
person university 48.00 45.00 52.00 43.00 47.00 46.00
plays pro sport 83.00 83.00 81.00 48.00 82.00 85.00
pokemon evolution 100.00 100.00 100.00 30.00 100.00 100.00
president birth year 0.00 0.00 0.00 0.00 0.00 0.00
president election year 0.00 0.00 0.00 0.00 0.00 0.00
product by company 85.00 82.00 85.00 38.00 85.00 84.00
star constellation name 87.00 87.00 87.00 75.00 87.00 87.00
superhero archnemesis 48.00 43.00 46.00 4.00 48.00 44.00
superhero person 76.00 73.00 75.00 5.00 77.00 72.00
AVG 73.08 70.33 73.06 25.20 73.03 70.61
Table 5: Per-relation knowledge recall accuracy results (%) on the LRE dataset for Qwen2.5-14B across different quantization methods.